OpenStudy (anonymous):

The sum of two consecutive odd integers is 80. What is the smaller number? A: 41 B: 43 C: 37 D: 39

OpenStudy (godlovesme):

consecutive #s are represented as x and x+1 so u will do x+(x+1)= 80 make sense so far? :)

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

makes sense

OpenStudy (asnaseer):

@Godlovesme - note that the question says "consecutive ODD integers"

OpenStudy (asnaseer):

e.g. 5 and 7 are consecutive ODD integers but 7 is NOT 5+1

OpenStudy (godlovesme):

im soo sorry =0 its x and x+2 right? @asnaseer D:

OpenStudy (asnaseer):

yup! :)

OpenStudy (anonymous):

is it C? @asnaseer

OpenStudy (godlovesme):

pheew lol anyways sorry guys ;0 x+ (x+2)=80 2x+2-2=80-2 2x/2=78/2 x=39 --> smaller #

OpenStudy (anonymous):

THANK YOU @Godlovesme <3

OpenStudy (godlovesme):

anytime ^.^

OpenStudy (anonymous):

Can I ask another?

OpenStudy (godlovesme):

sure

OpenStudy (solomonzelman):

I would use a different notation: I would make the smaller number x-1 and the larger number x+1 so I would be getting (x-1)+(x+1)=80, then the 1s cancel and it is faster a bit, this way,.

OpenStudy (anonymous):

If A = {3, 5, 7, 9} and B = {1, 2, 4, 6, 7, 9}, Find the intersection A: {1, 2, 3, 4, 5, 6, 7, 9} B: {1, 3, 5} C: {2, 4} D: {7, 9}

OpenStudy (solomonzelman):

which numbers do both sets have in common?

OpenStudy (anonymous):

Ohh okay now i get it sorry i forgot

OpenStudy (godlovesme):

yea find the common numbers in both sets

OpenStudy (anonymous):

A?

OpenStudy (anonymous):

or B

OpenStudy (solomonzelman):

okay, look you have set A with numbers 3, 5, 7, 9 and you have a set B with numbers 1, 2, 4, 6, 7, 9 can you find the number(s) that is part of Set A, and part of set B at once?

OpenStudy (solomonzelman):

and please don't just try to guess the answer, okay?

OpenStudy (godlovesme):

umm...  does that make sense?